Find the slope and y-intercept for the linear equation 36x + 6y = 24
Answer by checkley71(8403)

36X+6Y=24 OR 6Y=-36X+24 OR Y=-6X+4 THUS THE SLOPE=-6 & THE Y INTERCEPT=4